# Xiangyu(Tony) Zhang, Fei Tian, Xuerui Yang, Xiangyu Zhang, Daxin Jiang, Gang Yu

### v1

@@ Abstract
We present Step-Audio-EditX, the first open-source LLM-based audio model excelling at expressive and iterative audio editing—encompassing emotion, speaking style, and paralinguistics—alongside robust zero-shot text-to-speech (TTS) capabilities.Our core innovation lies in leveraging only large-margin synthetic data, which circumvents the need for embedding-based priors or auxiliary modules. This large-margin learning approach enables both iterative control and high expressivity across voices, and represents a fundamental pivot from the conventional focus on representation-level disentanglement. Evaluation results demonstrate that Step-Audio-EditX surpasses both MiniMax-2.6-hd and Doubao-Seed-TTS-2.0 in emotion editing and other fine-grained control tasks. Our code and models are available at https://github.com/stepfun-ai/Step-Audio-EditX.
@@ 5.1 Evaluation Benchmark
We introduce Step-Audio-Edit-Test, a benchmark that leverages LLM-as-a-judge model to evaluate model performance on emotion, speaking style, and paralinguistics. All evaluation audio is generated via zero-shot voice cloning and subsequently scored using the Gemini-2.5-Pro [^1] https://blog.google/products/gemini/gemini-2-5-pro-latest-preview/ model.
@@ 5.2 Evaluation Results
This section details our model’s performance on the Step-Audio-Edit-Test, benchmark and demonstrates its superior editing accuracy and scalability when used to edit audio generated by various closed-source TTS systems.
@@ 5.2.1 Emotion and Speaking Style Editing Results
Emotion $\uparrow$ Style $\uparrow$
@@
Chinese 57.0 71.7 74.5 77.7 41.6 62.1 65.8 69.2
English 49.9 60.5 61.5 63.7 50.3 62.4 64.3 63.1
Average 53.5 66.1 68.0 70.7 46.0 62.3 65.1 66.2
Chinese (Prompt-Fixed) 57.0 72.8 73.6 74.0 41.6 61.8 62.4 62.6
English (Prompt-Fixed) 49.9 60.4 61.3 61.3 50.3 65.1 63.1 64.0
Average 53.5 66.6 67.5 67.7 46.0 63.5 62.8 63.3
@@
Emotion $\uparrow$ Style $\uparrow$
@@
Chinese MiniMax-2.6-hd 71.6 78.4 77.6 82.4 37.9 58.6 64.0 66.9
Doubao-Seed-TTS-2.0 67.4 77.8 80.4 81.8 39.6 60.7 68.0 67.1
GPT-4o-mini-TTS 62.6 75.6 79.2 81.4 47.2 66.0 67.8 71.8
ElevenLabs-v2 60.4 70.8 79.0 76.8 43.1 61.8 69.4 70.9
English MiniMax-2.6-hd 55.0 64.8 64.6 66.0 50.9 62.1 63.1 64.3
Doubao-Seed-TTS-2.0 53.8 66.0 69.0 65.2 49.0 60.4 61.3 63.6
GPT-4o-mini-TTS 56.8 61.6 63.2 62.4 53.6 63.3 62.9 62.4
ElevenLabs-v2 51.0 61.6 64.0 64.0 49.0 63.0 63.7 62.0
Average MiniMax-2.6-hd 63.3 71.6 71.1 74.2 44.4 60.4 63.6 65.6
Doubao-Seed-TTS-2.0 60.6 71.9 74.7 73.5 44.3 60.6 64.7 65.4
GPT-4o-mini-TTS 59.7 68.6 71.2 71.9 50.4 64.7 65.4 67.1
ElevenLabs-v2 55.7 66.2 71.5 70.4 46.1 62.4 66.6 66.5
@@
Emotion Control on Closed-Source Models. Due to the limited availability of closed-source systems with emotion and speaking style control, this section presents a comparative evaluation of Doubao-Seed-TTS-2.0 and MiniMax-speech-2.6-hd, selected for their capability in both zero-shot cloning and emotion control. To meet the minimum audio length constraints of the closed-source models and to ensure a fair evaluation, the prompt audios for all the speakers in the Step-Audio-Edit-Test were extended in duration. These extended audios were employed for zero-shot cloning followed by two emotion editing iterations. Additionally, the cloned voices were used to generate emotional speech via each closed-source model’s native emotion control.The outputs from this native emotion control subsequently underwent one round of editing with our model. It can be observed from Table 3 that:
@@
Chinese Step-Audio-EditX 57.4 71.7 76.1 75.6
MiniMax-2.6-hd (Clone) 49.4 70.9 74.2 75.9
MiniMax-2.6-hd (Emotion Control) - 59.9 73.2 76.5
Doubao-Seed-TTS-2.0 (Clone) 50.8 68.8 75.0 77.2
Doubao-Seed-TTS-2.0 (Emotion Control) - 51.8 71.0 72.9
English Step-Audio-EditX 48.3 60.4 64.5 64.6
MiniMax-2.6-hd (Clone) 50.6 60.0 61.8 63.2
MiniMax-2.6-hd (Emotion Control) - 54.0 59.5 62.0
Doubao-Seed-TTS-2.0 (Clone) 47.1 58.3 60.9 62.1
Doubao-Seed-TTS-2.0 (Emotion Control) - 47.9 58.4 61.4
Average Step-Audio-EditX 52.9 66.1 70.3 70.1
MiniMax-2.6-hd (Clone) 50.0 65.5 68.8 69.6
MiniMax-2.6-hd (Emotion Control) - 57.0 66.4 69.3
Doubao-Seed-TTS-2.0 (Clone) 49.0 63.6 68.0 69.7
Doubao-Seed-TTS-2.0 (Emotion Control) - 49.9 64.7 67.2
@@ 5.2.2 Paralinguistic Results
Chinese MiniMax-speech-2.6-hd 1.73 2.80 2.88
Doubao-Seed-TTS-2.0 1.67 2.81 2.88
GPT-4o-mini-TTS 1.71 2.88 2.86
ElevenLabs-v2 1.70 2.71 2.87
English MiniMax-speech-2.6-hd 1.72 2.87 2.86
Doubao-Seed-TTS-2.0 1.72 2.75 2.90
GPT-4o-mini-TTS 1.90 2.90 2.85
ElevenLabs-v2 1.93 2.87 2.83
Average MiniMax-speech-2.6-hd 1.73 2.84 2.87
Doubao-Seed-TTS-2.0 1.70 2.78 2.89
GPT-4o-mini-TTS 1.81 2.89 2.86
ElevenLabs-v2 1.82 2.79 2.85

### v2

@@ Abstract
We present Step-Audio-EditX, the first open-source LLM-based reinforcement learning audio model excelling at expressive and iterative audio editing—encompassing emotion, speaking style, and paralinguistics—alongside robust zero-shot text-to-speech (TTS) capabilities. Our core innovation lies in leveraging only large-margin synthetic data in post-training, which circumvents the need for embedding-based priors or auxiliary modules. This large-margin learning approach enables both iterative control and high expressivity across voices, and represents a fundamental pivot from the conventional focus on representation-level disentanglement. Evaluation results demonstrate that Step-Audio-EditX surpasses both MiniMax-2.6-hd and Doubao-Seed-TTS-2.0 in emotion editing and other fine-grained control tasks. Our code and models are available at https://github.com/stepfun-ai/Step-Audio-EditX.
@@ 5.1 Evaluation Benchmark
We introduce Step-Audio-Edit-Benchmark, a benchmark that leverages LLM-as-a-judge model to evaluate model performance on emotion, speaking style, and paralinguistics. All evaluation audio is generated via zero-shot voice cloning and subsequently scored using the Gemini-2.5-Pro [^1] https://blog.google/products/gemini/gemini-2-5-pro-latest-preview/ model.
@@ 5.2 Evaluation Results
This section details our model’s performance on the Step-Audio-Edit-Benchmark, benchmark and demonstrates its superior editing accuracy and scalability when used to edit audio generated by various closed-source TTS systems.
@@ 5.2.1 Emotion and Speaking Style Editing Results
Emotion $\uparrow$ Speaking Style $\uparrow$
@@
Chinese 58.7 73.6 75.1 77.8 40.4 62.1 65.3 68.0
English 51.2 60.0 63.1 64.2 48.8 63.4 62.3 64.4
Average 55.0 66.8 69.1 71.0 44.6 62.8 63.8 66.2
Chinese (Prompt-Fixed) 57.5 73.1 76.3 75.8 41.1 62.0 65.1 63.7
English (Prompt-Fixed) 49.7 60.4 61.1 62.8 50.0 63.4 63.2 63.9
Average 53.6 66.8 68.7 69.3 45.6 62.7 64.2 63.8
@@
Emotion $\uparrow$ Speaking Style $\uparrow$
@@
Chinese MiniMax-2.6-hd 71.6 78.6 81.2 83.4 36.7 58.8 63.1 67.3
Doubao-Seed-TTS-2.0 67.4 77.8 80.6 82.8 38.2 60.2 65.0 64.9
GPT-4o-mini-TTS 62.6 76.0 77.0 81.8 45.9 64.0 65.7 69.7
ElevenLabs-v2 60.4 74.6 77.4 79.2 43.8 63.3 69.7 70.8
English MiniMax-2.6-hd 55.0 64.0 64.2 66.4 51.9 60.3 62.3 64.3
Doubao-Seed-TTS-2.0 53.8 65.8 65.8 66.2 47.0 62.0 62.7 62.3
GPT-4o-mini-TTS 56.8 61.4 64.8 65.2 52.3 62.3 62.4 63.4
ElevenLabs-v2 51.0 61.2 64.0 65.2 51.0 62.1 62.6 64.0
Average MiniMax-2.6-hd 63.3 71.3 72.7 74.9 44.2 59.6 62.7 65.8
Doubao-Seed-TTS-2.0 60.6 71.8 73.2 74.5 42.6 61.1 63.9 63.6
GPT-4o-mini-TTS 59.7 68.7 70.9 73.5 49.1 63.2 64.1 66.6
ElevenLabs-v2 55.7 67.9 70.7 72.2 47.4 62.7 66.1 67.4
@@
Emotion Control on Closed-Source Models. Due to the limited availability of closed-source systems with emotion and speaking style control, this section presents a comparative evaluation of Doubao-Seed-TTS-2.0 and MiniMax-speech-2.6-hd, selected for their capability in both zero-shot cloning and emotion control. To meet the minimum audio length constraints of the closed-source models and to ensure a fair evaluation, the prompt audios for all the speakers in the Step-Audio-Edit-Benchmark were extended in duration. These extended audios were employed for zero-shot cloning followed by two emotion editing iterations. Additionally, the cloned voices were used to generate emotional speech via each closed-source model’s native emotion control.The outputs from this native emotion control subsequently underwent one round of editing with our model. It can be observed from Table 3 that:
@@
Chinese Step-Audio-EditX 58.6 72.1 75.7 77.8
MiniMax-2.6-hd (Clone) 49.4 72.1 75.6 78.1
MiniMax-2.6-hd (Emotion Control) - 59.9 75.2 78.2
Doubao-Seed-TTS-2.0 (Clone) 50.8 70.9 75.6 76.4
Doubao-Seed-TTS-2.0 (Emotion Control) - 51.8 68.9 75.9
English Step-Audio-EditX 51.0 61.0 63.2 64.3
MiniMax-2.6-hd (Clone) 50.6 60.2 62.9 63.6
MiniMax-2.6-hd (Emotion Control) - 54.0 61.5 62.1
Doubao-Seed-TTS-2.0 (Clone) 47.1 57.7 61.9 64.5
Doubao-Seed-TTS-2.0 (Emotion Control) - 47.9 60.7 61.7
Average Step-Audio-EditX 54.8 66.6 69.5 71.1
MiniMax-2.6-hd (Clone) 50.0 66.2 69.3 70.9
MiniMax-2.6-hd (Emotion Control) - 57.0 68.4 70.2
Doubao-Seed-TTS-2.0 (Clone) 49.0 64.3 68.8 70.5
Doubao-Seed-TTS-2.0 (Emotion Control) - 49.9 64.8 68.8
@@ 5.2.2 Paralinguistic Results
Chinese MiniMax-speech-2.6-hd 1.73 2.80 2.90
Doubao-Seed-TTS-2.0 1.67 2.81 2.90
GPT-4o-mini-TTS 1.71 2.88 2.93
ElevenLabs-v2 1.70 2.71 2.92
English MiniMax-speech-2.6-hd 1.72 2.87 2.88
Doubao-Seed-TTS-2.0 1.72 2.75 2.92
GPT-4o-mini-TTS 1.90 2.90 2.88
ElevenLabs-v2 1.93 2.87 2.88
Average MiniMax-speech-2.6-hd 1.73 2.84 2.89
Doubao-Seed-TTS-2.0 1.70 2.78 2.91
GPT-4o-mini-TTS 1.81 2.89 2.90
ElevenLabs-v2 1.82 2.79 2.90